Feedback hi See Periodic Table See Hint Exactly 10.1 mL of water at 31.0°C are added to a hot iron skillet. All of the water is converted into steam at 100.0°C, The mass of the 3 104 pan is 1.50 kg and the molar heat capacity of iron is 25.19 J/(mol°C). What is the temperature change of the skillet? 104
Answer #1

Heat lose of iron           = heat gain of water

mc\Deltat                           = mc\Deltat

1500*0.45*(t-100)         = 10.1*4.18*(100-31)

     t   = 104.30C

temperature change of the skillet = 104.3-100 = 4.30c